Sarah has some stocks invested in the market. Last year it was worth$24,638. This year the stocks are worth $13,874. Is this an increase or a decrease? Write an equation to solve how to change the percentage change, rounded to the nearest tenth.
1
Expert's answer
2012-08-17T07:48:15-0400
C = 13874
C0 = 24638
C - C0 = 13874 - 24638 = -10764 < 0, so this is a
decrease.
For the percentage change P:
C = C0 * (100 + P) / 100
P = 100
* C / C0 - 100
P = -43.7%